OpenStudy (anonymous):

Help with conjugates (Algebra 2B) equation and explanation in comments.

OpenStudy (anonymous):

So I need to rationalize the denominator and simplify but I don't know how I would go about that since I can't figure out the conjugate. \[\frac{ 15\sqrt{6} }{ 3\sqrt{5} }\]

OpenStudy (solomonzelman):

multiply times √5 on top and bottom

OpenStudy (solomonzelman):